OpenStudy (anonymous):

y=2x^2-x-3 Please help! Solve for the real roots of the equation and explain how you would use the graph of the equation to check the values for the real roots.

myininaya (myininaya):

\[2x^2-x-3=0\] You need to find two factors of a(c) that have sum b You need to find two factors of 2(-3) that have sum -1 So 2(-3) is 2(-3) lol and 2+(-3)=-1 So boom Now we know 2x+(-3x)=-1x So we are going to replace middle term with 2x-3x or -3x+2x either one is fine they mean the same thing \[2x^2-3x+2x-3=0\] Now factor by grouping shall be done The first two terms have x in common so we can do \[x(2x-3)+2x-3=0\] The last two terms have 1 in common so we can do \[x(2x-3)+1(2x-3)=0\] So now we have two terms and they both have the factor (2x-3) in common So we can factor out (2x-3) like so \[(2x-3)(x+1)=0\]

myininaya (myininaya):

Setting both factors =0 and solving for x will give us the two real solutions :)

myininaya (myininaya):

\[2x-3=0 \text{ or } x+1=0\]

OpenStudy (anonymous):

Wow, thank you so much myininaya!

myininaya (myininaya):

Now when you you finish getting these answers algebraically you can compare them to the answers by finding them graphically just by seeing where the graph crosses the x-axis (the horizontal axis)

OpenStudy (anonymous):

Great! Thanks!

myininaya (myininaya):

Np :)